# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

### v2

@@ 2.2 Algorithm
In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1. KL-PCB first leverages least-square estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset. In Zhao et al. (2024), such ${\bar{g}}$ is directly applied to construct the estimated policy. In contrast, we construct a pessimistic estimator of ${g^{*}}$ following the well-known pessimism principle in offline RL (Jin et al., 2021). Specifically, we define the bonus term $\Gamma_{n}$ through the confidence radius $\beta=\sqrt{128\log\big{(}2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}$ as
@@
We then obtain our pessimistic estimation ${\widehat{g}}$ by setting ${\widehat{g}}={\bar{g}}-\Gamma_{n}$, which is less than ${g^{*}}$ with high probability. After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
@@
1:  Compute the least square estimation of reward function ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
@@ 2.3 Theoretical Results
Previously, Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.6. As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.7, which is weaker than Assumption 2.6. We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
@@
Previously, Zhao et al. (2024) provided a sample complexity lower bound of $\Omega(\eta\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)/{\epsilon})$ under KL-regularized objective. Foster et al. (2025) also provided a lower bound of $\Omega(C^{\pi^{*}})$ for KL-regularized objective to show the necessity of coverage. Compared to their results, our result shows that the multiplicative dependency on $C^{\pi^{*}}$ is necessary for the first time.
@@
Theorem 2.10 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\Omega(\eta C^{\pi^{*}})\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ samples to output an $\epsilon$-optimal policy. The presence of $\exp(\mathrm{poly}(\eta))$ in the range of $C^{*}$ is inevitable, since we always have $C^{\pi^{*}}\leq\exp(\eta)$ in reverse KL regularized bandits with bounded rewards. As discussed in Remark 2.8, when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ are constant, $D^{2}_{\pi^{*}}\lesssim C^{\pi^{*}}$, indicating that KL-PCB is near-optimal under this special case.
2.4 Proof of Theorem 2.9
In this section, we finish the proof of Theorem 2.9. At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t. ${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2025, Exercise I.37), and thus $J(\pi)$ is strongly concave. Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer. This intuition guides our analysis for contextual bandits.
We begin the proof with the definition of the event $\mathcal{E}(\delta)$ given $\delta>0$ as
where $\Gamma_{n}$ is defined in (2.3). Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{g}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function ${g^{*}}$. More specifically, we have the following lemma, whose proof are deferred to Appendix B.1.
@@
The following covariance-type observation is the first pivot.
@@
If a bounded random variable $X\leq 0$ almost surely, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
Remark 2.14.
While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived. The intuition behind Lemma 2.13 is very natural: $X$ and $X^{2}$ cannot be positively correlated.
We further define the following quantities. For all $\gamma\in[0,1]$, we define $g_{\gamma}\coloneqq\gamma{\widehat{g}}+(1-\gamma){g^{*}}$ and further denote
Lemma 2.15.
Proof.
$\ \leq 0,$
where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.13. ∎
Finally, we need the following lemma to bound the performance difference between two policy w.r.t. the KL-regularized objective $J$. For any given $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, we define $\pi_{g}(\cdot|s)\propto\exp\big{(}\eta\cdot g(s,\cdot)\big{)},\forall s\in{\mathcal{S}}$ to facilitate presentation; then $\pi^{*}=\pi_{{g^{*}}}$ by definition. Please refer to Appendix B.3 for the proof of the lemma.
Lemma 2.16.
Given $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, $\exists\gamma\in[0,1]$ such that for $g_{\gamma}=\gamma g+(1-\gamma)g^{*}$ and $\pi_{\gamma}\coloneqq\pi_{g_{\gamma}}$,
$J(\pi^{*})-J(\pi_{g})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big{[}\big{(}{g^{*}}-g\big{)}^{2}(s,a)\big{]}.$
Now we are ready to prove Theorem 2.9.
Proof of Theorem 2.9.
$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (2.5)
where the first inequality holds due to Lemma 2.16 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.15. The term $G(0)$ can be further bounded with the $D^{2}$-based concentrability as follows
$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}\log_{{\mathcal{G}}}(\epsilon_{c})),$ (2.6)
where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma 2.15, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2. By Lemma 2.12, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof. ∎
@@ 3 $f$-divergence-regularized Contextual Bandits
As we discussed in Section 2, the fast rate is primarily achieved due to the strong convexity of $\pi\mapsto\mathsf{KL}(\cdot\|\pi^{\mathsf{ref}})$. However, KL is just an instance of $f$-divergence with $f(x)=x\log x$, which is not strongly convex, viz., only locally strongly convex. Motivated by this observation, we examine the use of $f$-divergence regularization with strongly convex $f$, which, in principle, may introduce a more favorable curvature in the performance metric of offline learning.
@@ 3.1 Problem Setup
We study the same contextual bandit problems which follows the definition in Section 2.1. In this section, we consider the following $f$-divergence regularized objective
$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}[r(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]},$ (3.1)
where $\eta$ is the inverse temperature regularization intensity and $D_{f}(\pi||\pi^{\mathsf{ref}})$ is given by
$D_{f}(\pi||\pi^{\mathsf{ref}})=\mathbb{E}_{a\sim\pi^{\mathsf{ref}}}\bigg{[}f\bigg{(}\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{)}\bigg{]}.$
We consider those functions $f:(0,+\infty)\to\mathbb{R}$ with nice positive curvature as follows.
@@
Many elementary functions like quadratic polynomials naturally satisfy Assumption 3.1. For instance, the 1-strongly convex $f(x)=(x-1)^{2}/2$ yields $D_{f}(P\|Q)=\chi^{2}(P\|Q)$, which is the popular $\chi^{2}$-divergence recently considered in RL literature (see e.g., Zhan et al. (2022); Huang et al. (2024); Amortila et al. (2024)) and exhibits a promising theoretical potential on relaxing the data coverage requirement for efficient offline policy learning, at least against the unregularized objective.
@@ 3.2 Algorithm and Main Results
1:  Compute the least square estimation of reward function
${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
@@
The following hardness result justify the near-optimality of Theorem 3.2 for $f$-divergence-regularized bandits.
@@
For any $\epsilon\in(0,1),\alpha>0,\eta>0$, and algorithm Alg, there is an $\alpha$-strongly-convex function $f$ and an $f$-divergence-regularized contextual bandit instance such that Alg requires at least $\Omega\big{(}\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big{)}$ samples to return an $\epsilon$-optimal policy.
3.3 Insights Behind the Proof of Theorem 3.2
We provide an overview of key analysis techniques for proving Theorem 3.2. For the simplicity of presentation, we consider multi-arm bandits here and omit the subscript for context $s$.
We consider the function $H(\pi)=\eta^{-1}D_{f}(\pi\|\pi^{\mathsf{ref}})$, which is the regularizer in the objective. Then its convex conjugate is given by $H^{*}(r)=\sup_{\pi\in\Delta^{d}}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$, which is exactly the expected reward obtained by the optimal policy given reward function $r$. One observation is that when $f$ is strongly convex, the induced $f$-divergence, and therefore the function $H$ is also strongly convex. Therefore, let $\pi_{r}=\mathop{\mathrm{argmax}}_{\pi}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$ given some reward function $r$, the strong convexity of $H(\pi)$ gives that $\nabla H^{*}(r)=\pi_{r}$. This leads to the following regret decomposition, which is one of our key observations.
$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(a)]$
$\quad-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}$
@@
which is the Bregman divergence of the dual function $H^{*}$ and therefore can be bounded by $({g^{*}}-{\bar{g}})^{\top}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ for some $\widetilde{g}$. By Proposition 3.2 in Penot (1994), we can bound $H^{*}(\widetilde{g})$ as follows
$H^{*}(\widetilde{g})$ $\preceq\big{(}\nabla^{2}H(\pi_{\widetilde{g}})\big{)}^{-1}$
$\preceq\alpha^{-1}\eta\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|})\big{)},$

### v3

@@ 2.2 Algorithm
In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1. KL-PCB first leverages least-square estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset. In Zhao et al. (2024), such $\bar{g}$} is directly applied to construct the estimated policy. In contrast, we construct a pessimistic estimator of $g^{*}$} following the well-known pessimism principle in offline RL (Jin et al., 2021). Specifically, we define the bonus term $\Gamma_{n}$ through the confidence radius $\beta=\sqrt{128\log\big(2\mathcal{N}_{{\mathcal{G}}}(\epsilon)/\delta\big)/3n+18\epsilon}$ as
@@
We then obtain our pessimistic estimation $^{\widehat{g}}$ by setting ${\widehat{g}}={\bar{g}}-\Gamma_{n}$, which is less than $g^{*}$} with high probability. Formally, let the event $\mathcal{E}(\delta)$ given $\delta>0$ defined as
on which the least square estimation $\bar{g}$} obtained in Line 1 of Algorithm 1 does not deviate too much from the true function $g^{*}$} and therefore $^{\widehat{g}}$ is a pessimistic estimation of $g^{*}$}. We have the following lemma indicating that this event holds with high probability.
Lemma 2.9.
After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
@@
1: Least square estimation of reward function ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big(g(s_{i},a_{i})-r_{i}\big)^{2}$
@@ 2.3 Theoretical Results
Previously, Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.7. As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.8, which is weaker than Assumption 2.7. We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem, which, together with Theorem 2.10, demonstrates that single-policy concentrability is both necessary and sufficient for near-optimal offline learning evaluated by KL-regularized objectives.
@@
Previously, Zhao et al. (2024) provided a sample complexity lower bound of $\Omega(\eta\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)/{\epsilon})$ under KL regularization. Foster et al. (2025) also provided a lower bound of $\Omega(C^{\pi^{*}})$ for KL-regularized objective to show the necessity of coverage. Compared to their results, our result shows that the multiplicative dependency on $C^{\pi^{*}}$ is necessary for the first time.
@@
Theorem 2.11 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\Omega(\eta C^{\pi^{*}})\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ samples to output an $\epsilon$-optimal policy. The presence of $\exp(\mathrm{poly}(\eta))$ in the range of $C^{*}$ is inevitable, since we always have $C^{\pi^{*}}\leq\exp(\eta)$ in reverse KL regularized bandits with bounded rewards.
Remark 2.13.
As discussed before, we might have some easy instances with $D^{2}_{\pi^{*}}\leq C^{\pi^{*}}$, where KL-PCB outperforms the lower bound. This does not volates Theorem 2.11 since Theorem 2.11 only guarantees that there exist some hard instances that all algorithms require at least $\Omega(\min\{\eta\epsilon^{-1},\epsilon^{-2}\}C^{*}\log\mathcal{N}_{G}(\epsilon))$ samples.
2.4 Proof Overview of Theorem 2.10
[…]
@@
Let $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ be any reward function, then there exist some $\gamma\in[0,1]$ such that the sub-optimality gap of $\pi_{g}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big(\eta g(s,\cdot)\big)$ can be bounded as
$J(\pi^{*})-J(\pi_{g})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big[\big({g^{*}}-g\big)^{2}(s,a)\big],$
where $g_{\gamma}\coloneqq\gamma g+(1-\gamma)g^{*}$ and $\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big(\eta g_{\gamma}(s,\cdot)\big)$.
In Zhao et al. (2024), because the $g$ in Lemma 2.14 is substituted with only the least-square estimator $\bar{g}$ with no extra structures, the reliance on the “mid-point” policy $\pi_{\gamma}$ can only be controlled all-policy concentrability. However, our $g$ is the pessimistic estimator $\widehat{g}$ of $g^{*}$} in Algorithm 1, and thus the presence of $\pi_{\gamma}$ can be eliminated for free: let $G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big[\big({\widehat{g}}-{{g^{*}}}\big)^{2}(s,a)\Big]$ and $\triangle(s,a)\coloneqq\big({\widehat{g}}-{{g^{*}}})(s,a)\leq 0$, then a direct computation (detailed in the proof of Lemma D.3) yields
$G^{\prime}(\gamma)=\eta\mathbb{E}_{\rho}\Big[\mathbb{E}_{\pi_{\gamma}}\big[\triangle^{3}(s,a)\big]-\mathbb{E}_{\pi_{\gamma}}\big[\triangle^{2}(s,a)\big]\mathbb{E}_{\pi_{\gamma}}\big[\triangle(s,a)\big]\Big]\leq 0.$ (2.5)
This gives $J(\pi^{*})-J(\widehat{\pi})\leq\eta\mathbb{E}_{\rho\times\pi^{*}}\big[({\widehat{g}}-{{g^{*}}})^{2}(s,a)\big]$, which can be bounded with single-policy concentrability while still achieves the sharp dependency $\epsilon^{-1}$ on $\epsilon$. Here, (2.5) holds due to a moment-based machinery in Lemma 2.15.
@@
If $\mathbb{P}(X\leq 0)=1$ and $\mathbb{E}|X|^{3}<\infty$, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
[…]
@@ 3 $f$-divergence-regularized Contextual Bandits
As discussed in Section 2, the fast rate implied by Theorems 2.10 and 2.11 is primarily achieved due to the strong convexity of $\pi\mapsto\mathsf{KL}(\pi\|\pi^{\mathsf{ref}})$. However, KL is just an instance of $f$-divergence with $f(x)=x\log x$, which is only locally strongly convex but not strongly convex. Motivated by this observation, we further examine $f$-divergence regularization with strongly convex $f$, which may introduce a more favorable curvature in the performance metric of offline learning in principle.
@@ 3.1 Problem Setup
We study a contextual bandit setting similar to that in Section 2.1. In this section, we consider the following $f$-divergence regularized objective
$J_{\eta,D_{f}}(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}[r(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big[D_{f}\big(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big)\big],$ (3.1)
where $\eta$ is the regularization intensity and $D_{f}(p\|q)\coloneqq\mathbb{E}_{a\sim q}\Big[f\big({p(a)}/{q(a)}\big)\Big]$ is the $f$-divergence. Let the optimal policy be $\pi^{*}_{\eta,D_{f}}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J_{\eta,D_{f}}(\pi)$ and we re-define the learning objective as searching for a policy $\pi$ with $\mathrm{SubOpt}_{f\mathrm{div}}(\pi)\coloneqq J(\pi^{*})-J(\pi)\leq\epsilon$. [^4] We again suppress $J_{\eta,D_{f}}(\cdot)$ into $J(\cdot)$ and $\pi^{*}_{\eta,D_{f}}$ into $\pi^{*}$ when there is no confusion. We consider those functions $f:(0,+\infty)\to\mathbb{R}$ with a nice positive curvature condition in Assumption 3.1.
@@
Many elementary functions like quadratic polynomials naturally satisfy Assumption 3.1. For instance, the 1-strongly convex $f(x)=(x-1)^{2}/2$ induces $D_{f}(P\|Q)=\chi^{2}(P\|Q)$, which is the $\chi^{2}$-divergence recently considered in RL literature (see e.g., Zhan et al. (2022); Huang et al. (2025b); Amortila et al. (2024)). This regularization exhibits a promising theoretical potential for relaxing the data coverage requirement for efficient offline policy learning (Huang et al., 2025b) and to be effective in preventing reward hacking (Laidlaw et al., 2025) against unregularized objectives. These favorable benefits are primary due to the observation that strongly convex $f$’s impose a stronger penalization on actions out of the coverage of $\pi^{\mathsf{ref}}$.
@@ 3.2 Algorithm and Main Results
1: Least square estimation ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big(g(s_{i},a_{i})-r_{i}\big)^{2}$
@@
The following hardness result justify the near-optimality of Theorem 3.2 for $f$-divergence-regularized contextual bandits.
@@
For any $\epsilon\in(0,1),\alpha>0,\eta>0,S>32/3\cdot\log 2$, sufficiently small $\epsilon$, and algorithm Alg, there is an $\alpha$-strongly-convex function $f$ and an $f$-divergence-regularized contextual bandit instance such that Alg requires at least $\Omega\big(\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big)$ samples to return an $\epsilon$-optimal policy.
3.3 Proof Overview of Theorem 3.2
We provide an overview of key analysis techniques for proving Theorem 3.2. Unlike KL-regularization, the $\pi^{*}$ under $f$-divergence might not have a closed form. This means that the proof of Lemma 2.14, which relies on the closed form of $\pi^{*}$, cannot be directly adopted. Therefore, we address this from a dual-Bregman perspective, inspired by Abernethy et al. (2015). For the simplicity of presentation, we consider multi-armed bandits here and omit the subscript for context $s$.
We consider the function $H(\pi)=\eta^{-1}D_{f}(\pi\|\pi^{\mathsf{ref}})$, which is the regularizer in the objective. Then its convex conjugate is given by $H^{*}(r)=\sup_{\pi\in\Delta^{d}}\{\left\langle\pi,r\right\rangle-H(\pi)\}$, which is exactly the expected reward obtained by the optimal policy given reward function $r$. One observation is that when $f$ is strongly convex, the induced $f$-divergence, and therefore the function $H$ are also strongly convex. Therefore, let $\pi_{r}=\mathop{\mathrm{argmax}}_{\pi}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$ given some reward function $r$, the strong convexity of $H(\pi)$ gives that $\nabla H^{*}(r)=\pi_{r}$. This leads to the following regret decomposition, which is one of our key observations:
[…]
@@
[…]